# Time-domain identification of distinct mechanisms for competing charge density waves in a rare-earth tritelluride

### v1

@@ Abstract
Understanding phase transitions and phase competition in solid-state systems is a central pursuit in condensed matter physics and materials science. In modern statistical mechanics, phase transitions are generally classified as either first-order or second-order, based on the continuity properties of the free energy and its derivatives1. Despite decades of research, the experimental classification and underlying mechanisms of phase transitions in many physical systems remain elusive. A major challenge lies in the difficulty of directly probing the free energy landscape in equilibrium, which governs these transitions. An alternative approach is to drive the system out of equilibrium using an ultrafast laser pulse that transiently weakens or quenches the ordered phase. As the system relaxes back to equilibrium along with its free energy surface, tracking the temporal evolution of relevant observables can provide insights into both the underlying free energy landscape and the mechanisms driving the phase transition2, 3, 4, 5, 6, 7, 8, 9.
@@
A prominent example of CDW with unidentified origin is those in rare-earth tritelluride (Te3, where denotes a rare-earth element), a family of prototypical CDW materials. Te3 possess a layered, quasi-tetragonal structure [Fig. 1(a)] with a slight in-plane anisotropy ( at 300 K)15 leading to a preferred direction of the CDW along the -axis. This dominant CDW is hereafter referred to as -CDW. The Fermi surface, which is similar for all Te3, arises from the Te bilayer square-net sheets 16. The normal-state Fermi surface calculated from the tight-binding model is depicted in the lower-right corner of (Fig. 1a). The rare-earth element can be used as a tuning knob to adjust the chemical pressure and thus the ground state of the material. Materials with light rare-earth atoms only undergo one CDW transition, forming a -CDW below the transition temperature (). In materials with rare-earth atoms heavier than Gd, a second CDW order, typically referred to as -CDW, emerges along the -axis, perpendicular to the -axis in the same Te sheet, at a lower transition temperature ()(See Fig. 1b, detailed discussion on the nomenclature can be found in Supplemental Information). Though sharing almost the same structural motifs and nesting vectors on the Fermi surface, the -CDW is not simply the -CDW rotated by 90∘. One prominent difference is that the -CDW is not associated with a soft phonon mode around from inelastic X-ray measurements17 (see Fig. 1b), arguing against a stereotypical second-order CDW phase transition 10, 18. On the other hand, a recent Raman scattering study on Te3 observed amplitude mode softening towards transition temperature (ref. 19), implicitly contradicting the previous X-ray scattering and optical measurements 17, 20, 21, 22, 23. Moreover, recent transport measurements discovered a strong violation of Wiedemann-Franz law below , indicating the breakdown of typical electron-phonon coupling picture of CDW formation 24. These findings point toward an exotic origin of the -CDW that motivates the current study.
@@
The observation that the recovery time of the -CDW is nearly independent of the quench strength can be understood in an intuitive physical picture. Upon photoexcitation, some domains of the -CDW are suppressed. Each domain can be defined with respect to a defect in the center of the domain. These defects function as seeds for the nucleation of the -CDW. Upon cooling of the electronic subsystem, a new minimum with lower energy forms. The height of the local free energy barrier determines the probability of nucleating into the new minima. Once the local nucleation happens, the mean-field value is suddenly large and fluctuations become negligible. The bubble-like CDW domain then grows spherically like an inflating bubble at the rate of nucleation until it impinges another growing bubble (see Fig. 4c and Supplemental Information Sec. II.B.2). The recovery rate of the -CDW is then governed by the nucleation rate set by the pressure difference, arising from the difference in free energy, between the -CDW phase and the phase without the a-CDW order, independent of light-induced fluctuations. An increase in laser fluence or, equivalently, the quench strength increases the number of nucleation bubbles but not the recovery rates. The recovery time of -CDW is thus barely fluence-dependent.
The TDGL theory and the physical picture described above should, in principle, be generally applicable to other systems beyond ErTe3. Comparing previous pump-probe experiments that measured the order parameter amplitude of some prototypical CDWs, as shown in Fig. 4d, we find that in material systems with first-order CDW phase transitions, such as 1-TaS2, IrTe2, and indium nanowires on silicon substrates 32, 33, 34, the recovery timescales are insensitive to the laser fluence in the fluence regime we studied in this work. While in systems with second-order CDW transitions, such as LaTe3, 1-TiSe2, and K0.3MoO3 (blue bronze) 4, 35, 36, the recovery slows down with a growing fluence, due to the light-induced fluctuation (see Sec. III in Supplemental Information for details). From this perspective, ErTe3 provides a unique platform as a controlled comparison between CDW transitions of different origins, as marked in Fig. 4d with red and blue diamonds. However, we notice that exceptions to this simple time-domain classification may occur in some transient reflectivity measurements where the time evolution of reflectivity involve significant contributions from other physical processes 49. This, on the other hand, also demonstrates the advantage of trARPES measurements in isolating the dynamics of the order parameter. This comparative study, though not exhaustive, supports the validity of our time-domain protocol based on TDGL theory.
@@ I.1 Material preparation
Single crystals were grown via a self-flux technique. Elements in the molar ratio Te1-x, , were put into alumina crucibles and vacuum sealed in quartz tubes. The mixtures were heated to 800–900∘C and slowly cooled to end temperatures in the range of 500–600∘C. The remaining melt was decanted in a centrifuge. Before trARPES measurements, bulk single crystals were cleaved in ultrahigh vacuum ( Torr) to expose a pristine surface.

### v2

@@ Abstract
In modern statistical mechanics, phase transitions are generally classified as either first-order or second-order, based on the continuity properties of the free energy and its derivatives1. Despite decades of research, the experimental classification and underlying mechanisms of phase transitions in many physical systems remain elusive. A major challenge lies in the difficulty of directly probing the free energy landscape in equilibrium, which governs these transitions. An alternative approach is to drive the system out of equilibrium using an ultrafast laser pulse that transiently weakens or quenches the ordered phase. As the system relaxes back to equilibrium along with its free energy surface, tracking the temporal evolution of relevant observables can provide insights into both the underlying free energy landscape and the mechanisms driving the phase transition2, 3, 4, 5, 6, 7, 8, 9.
@@
A prominent example of CDW with unidentified origin is those in rare-earth tritelluride (Te3, where denotes a rare-earth element), a family of prototypical CDW materials. Te3 possess a layered, quasi-tetragonal structure [Fig. 1a] with a slight in-plane anisotropy ( at 300 K)15 leading to a preferred direction of the CDW along the -axis. This dominant CDW is hereafter referred to as -CDW. The Fermi surface, which is similar for all Te3, arises from the Te bilayer square-net sheets 16. The normal-state Fermi surface calculated from the tight-binding model is depicted in the lower-right corner of (Fig. 1a). The rare-earth element can be used as a tuning knob to adjust the chemical pressure and thus the ground state of the material. Materials with light rare-earth atoms only undergo one CDW transition, forming a -CDW below the transition temperature (). In materials with rare-earth atoms heavier than Gd, a second CDW order, typically referred to as -CDW, emerges along the -axis, perpendicular to the -axis in the same Te sheet, at a lower transition temperature ()(See Fig. 1b, detailed discussion on the nomenclature can be found in Supplemental Information). Though sharing almost the same structural motifs and nesting vectors on the Fermi surface, the -CDW is not simply the -CDW rotated by 90∘. One prominent difference is that the -CDW is not associated with a soft phonon mode around from inelastic X-ray measurements17 (see Fig. 1b), arguing against a stereotypical second-order CDW phase transition 10, 18. On the other hand, a recent Raman scattering study on Te3 observed amplitude mode softening towards transition temperature (ref. 19), implicitly contradicting the previous X-ray scattering and optical measurements 17, 20, 21, 22, 23. Moreover, recent transport measurements discovered a strong violation of Wiedemann-Franz law below , indicating the breakdown of typical electron-phonon coupling picture of CDW formation 24. These findings point toward an exotic origin of the -CDW that motivates the current study.
@@
The observation that the recovery time of the -CDW is nearly independent of the quench strength can be understood in an intuitive physical picture. Upon photoexcitation, some domains of the -CDW are suppressed. Each domain can be defined with respect to a defect in the center of the domain. These defects function as seeds for the nucleation of the -CDW. Upon cooling of the electronic subsystem, a new minimum with lower energy forms. The height of the local free energy barrier determines the probability of nucleating into the new minima. Once the local nucleation happens, the mean-field value is suddenly large and fluctuations become negligible. The bubble-like CDW domain then grows spherically like an inflating bubble at the rate of nucleation until it impinges another growing bubble (see Fig. 4c and Supplemental Information Sec. II.B.2). The recovery rate of the -CDW is then governed by the nucleation rate set by the pressure difference, arising from the difference in free energy, between the -CDW phase and the phase without the -CDW order, independent of light-induced fluctuations. An increase in laser fluence or, equivalently, the quench strength increases the number of nucleation bubbles but not the recovery rates. The recovery time of -CDW is thus barely fluence-dependent.
The TDGL theory and the physical picture described above should, in principle, be generally applicable to other systems beyond ErTe3. Comparing previous pump-probe experiments that measured the order parameter amplitude of some prototypical CDWs, as shown in Fig. 4d, we find that in material systems with first-order CDW phase transitions, such as 1-TaS2, IrTe2, and indium nanowires on silicon substrates 32, 33, 34, the recovery timescales are insensitive to the laser fluence in the fluence regime we studied in this work. While in systems with second-order CDW transitions, such as LaTe3, 1-TiSe2, and K0.3MoO3 (blue bronze) 4, 35, 36, the recovery slows down with growing fluence due to light-induced fluctuations (see Sec. III in Supplemental Information for details). From this perspective, ErTe3 provides a unique platform as a controlled comparison between CDW transitions of different origins, as marked in Fig. 4d with red and blue diamonds. However, we notice that exceptions to this simple time-domain classification may occur in some transient reflectivity measurements where the time evolution of reflectivity involve significant contributions from other physical processes 49. This, on the other hand, also demonstrates the advantage of trARPES measurements in isolating the dynamics of the order parameter. This comparative study, though not exhaustive, supports the validity of our time-domain protocol based on TDGL theory.
@@ I.1 Material preparation
Single crystals were grown via a self-flux technique. Elements in the molar ratio Te1-x, –0.003, were put into alumina crucibles and vacuum sealed in quartz tubes. The mixtures were heated to 800–900∘C and slowly cooled to end temperatures in the range of 500–600∘C. The remaining melt was decanted in a centrifuge. Before trARPES measurements, bulk single crystals were cleaved in ultrahigh vacuum ( Torr) to expose a pristine surface.